OpenStudy (anonymous):

A right triangle has a 50 degree angle. The hypotenuse is 10 cm long. To the nearest tenth, what is the length of the side opposite the 50 degree angle?

OpenStudy (jdoe0001):

 recall, -> http://www.mathwarehouse.com/trigonometry/images/sohcohtoa/sohcahtoa-all.png thus \(\bf sin(\theta)=\cfrac{opposite}{hypotenuse}\implies hypotenuse\cdot sin(\theta)=opposite\)

OpenStudy (jdoe0001):

make sure your calculator is in Degrees mode

OpenStudy (anonymous):

Sorry I was walking my dog

OpenStudy (jdoe0001):

I see people often times with a dog walking somehow it looks to me the Dog is walking them though =)

OpenStudy (anonymous):

Yea lol